A tree grows 1/2 inch per year. The tree is currently 50 feet tall; how long will it take to grow to 60 feet? Answers must be shown as a one variable algebraic equation with 'x' equaling the same value on each side of the equation. I believe the correct answer is 240 years, but I don't know how to show the work in the proper way.
Answer by stanbon(75887)

A tree grows 1/2 inch per year. The tree is currently 50 feet tall; how long will it take to grow to 60 feet? Answers must be shown as a one variable algebraic equation with 'x' equaling the same value on each side of the equation. I believe the correct answer is 240 years, but I don't know how to show the work in the proper way.
-----------------------------------------------
Slope is always the amount y grows when x increases by one.
---
slope = 1/2 inch per year
-----
intercept is always the value of y at zero, or at the "beginning".
-------------
intercept = 50 feet = 600 inches
===================================================================
height = (1/2)t+600
-----
How long to get to a height of 60 ft. or 720 inches
---
(1/2)t + 600 = 720
(1/2)t = 120
t = 240 years
=======================
Cheers,
Stan H.